OpenStudy (zenmo):

Find the 2nd derivative. Parametric curves.

OpenStudy (zenmo):

Find.\[\frac{ d^2y }{ dx^2 }\] \[x=e^t, y=te^{-1}\]

OpenStudy (zenmo):

So, far I have: \[x' = e^t, y' = e^{-t}(1-t), \frac{ dy }{ dx }=(1-t)e^{-2t}\]

OpenStudy (zenmo):

\[\frac{ d^2y }{ dx^2 }=\frac{ \frac{ d }{ dt }*(y') }{ x' }\]

OpenStudy (zenmo):

what is d/dt?

OpenStudy (baru):

try writing dy/dx in terms of x and y

OpenStudy (zale101):

Did you solve for t in one of the two equation and plugged it to the other equation?

OpenStudy (zale101):

Like what you used to do in systems of equations.

OpenStudy (zale101):

You want to get rid of the t-variable here.

OpenStudy (zenmo):

for d/dt?

OpenStudy (zale101):

You want to get your equation to look like y=f(x). To do that you need to combine both equations x=x(t) and y=y(t)

OpenStudy (zale101):

I have no clue where the d/dt came from

OpenStudy (zale101):

Any ideas @wio ?

OpenStudy (anonymous):

I think that \[ \Large \frac{d^2y}{dx^2} = \frac{d\left(\frac{dy}{dx}\right)}{dx} = \frac{\frac{d\left(\frac{dy}{dx}\right)}{dt}}{\frac{dx}{dt}} \]

OpenStudy (anonymous):

We already know \[\Large \frac{dy}{dx} = \frac{\frac{dy}{dt}}{\frac{dx}{dt}} \]

OpenStudy (anonymous):

So, I think we just need to replace \(y\) with \(\frac{dy}{dx}\).

OpenStudy (zenmo):

The source of the d/dt equation

OpenStudy (anonymous):

Well, that is what I was explaining.

OpenStudy (zenmo):

yea but what do we plug in for d/dt?

OpenStudy (anonymous):

\[ \frac{d(\ldots)}{dt} \]Is the derivative of \(\ldots\) with respect to \(t\). Just like \(\frac{dy}{dt}\).

OpenStudy (anonymous):

It's not a value, it's an operation you perform.

OpenStudy (baru):

\[\frac{dy}{dx} = z = (1-t)e^{-2t}\] then \[\frac{dz}{dx}= \frac{dz}{dt} \times \frac{dt}{dx}\]

OpenStudy (anonymous):

Can you differentiate \(dy/dx\) with respect to \(t\)? That is what \(\frac{d(dy/dx)}{dt}\) means.

OpenStudy (zenmo):

Would I need to differentiate \[(1-t)e^{-2t}\] ?

OpenStudy (anonymous):

Yes

OpenStudy (baru):

\[\frac{d^2 y}{dx^2} = \frac{d}{dt}((1-t)e^{-2t}) \times (x')^{-1}\]

OpenStudy (zenmo):

How would I simplify \[(-1)e^{-2t}+(1-t)(-2e^{-2t})\]?

OpenStudy (baru):

re-write it again in terms of x and y \[x=e^{-t} => x^2=e^{-2t}\\x-y=(1-t)e^{-t}\]

OpenStudy (baru):

\[x(x-y)=(1-t)e^{-2t}\]

OpenStudy (baru):

use those

OpenStudy (anonymous):

Did you divide by \(x'\) yet?

OpenStudy (zenmo):

not yet, i am attempting to simplify as previously posted first

OpenStudy (anonymous):

Factor out the \(e^{-2t}\)

OpenStudy (zenmo):

Okay, got it, the 2nd derivative is \[(2t-3)e^{-3t}\]

OpenStudy (zenmo):

Thanks all :)